Suppose an individual invests $40,000 in a load mutual fund for two years. The load fee entails an up-front commission charge of 3 percent of the amount invested and is deducted from the original funds invested. In addition, annual fund operating expenses (or 12b-1 fees) are 0.60 percent. The annual fees are charged on the average net asset value invested in the fund and are recorded at the end of each year. Investments in the fund return 5 percent each year paid on the last day of the year. If the investor reinvests the annual returns paid on the investment, calculate the annual return on the mutual fund over the two-year investment period. (Do not round intermediate calculations. Round your answer to 3 decimal places. (e.g., 32.161))
Answer #1

Annualized load fee = 3%/2 = 1.5%

Annual fund operating expenses = 0.6%

Total annual cost = 1.5%+0.6% = 2.1%. Annual return = 5% - 2.1% = 2.9%

Annual return calculation based on present value of investment:

Initial investment = $40,000. Front end load of 3% = 3% of 40,000 = $1,200

Total investible funds = 40,000 – 1,200 = $38,800

Investment value at the end of 1 year = 1.05*38800 = $40,740

Operating expenses based on average NAV = (40740+38800)/2 * 0.6% = 238.62

Thus net investible funds for year 2 = 40740-238.62 = 40,501.38

Investment value at the end of year 2 = 40,501.38*1.05 = 42526.45

Operating expenses based on average NAV = (40501.38+42526.45)/2*0.6% = 249.08

Net investment at end of year 2 = 42526.45-249.08 = 42277.37

Average annual compound return : 42277.37 = 40000*(1+g)^2

Or (1+g)^2 = 1.05693

Or 1+ g = 1.02807

Or g = 2.807%